OpenStudy (anonymous):

a line goes through the ordered pairs (8,-3) and (-4,6). Write an equation in Point-slope form, slope intercept form and standard form

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (radar):

The slope is equal to :\[\Delta y \over \Delta x\] then slope is: \[6-(-3) \over -4-8\] slope or m = 9/-12=- 3/4 Using the formula for a line, y=mx+b. we can take your given point (-4,6) and plug in the values that we now have: 6=-3/4 (-4) +b solving for b, 6=3+b or b=3 the equation is y=-3/4 (x)+3 that is known as the slope intercept form.

OpenStudy (radar):

\[(y-y _{1})=-3/4(x-x _{1})\]This is the point slope form, just plug in the x,y from your favorite point.

OpenStudy (radar):

like (y-6)= -3/4(x-(-4))

OpenStudy (radar):

Standard form would be for your line -3/4 (x)-y+3=0 in the form of Ax+By+C=0

OpenStudy (radar):

Hope you follow this with comprehension and that it proved helpful.